OpenStudy (anonymous):

\[\int \frac{dt}{\sqrt(t^2 -6t +13)}\]

OpenStudy (anonymous):

is the idea to complete the square the radical?

OpenStudy (anonymous):

i was thinking trig substitution...but don't know how...(that's the section)

OpenStudy (anonymous):

you have one step before the trig sub write \(x^2-6x+13=(x-3)^2+4\) to get \[\int\frac{dt}{\sqrt{(t-3)^2+4}}\] and the call say \(u=t-3\) and finally a trig sub

OpenStudy (anonymous):

at this point use usual one for \(\sqrt{u^2+b^2}\) use \(u=b\tan(\theta)\)

OpenStudy (anonymous):

\[u=bsec(\theta)\] ?

OpenStudy (anonymous):

maybe it does not matter lets check

OpenStudy (anonymous):

\[1+tan^2(\theta)=sec^2(\theta)\] \[\sqrt(1+tan^2(\theta))=sec(\theta)\] b=2

OpenStudy (anonymous):

\[\sqrt{u^2+4}\] \(u=2\tan(\theta), du=\sec^2(\theta)d\theta\) \[\sqrt{(2\tan^2(\theta)+4}=\sqrt{4(tan^2(\theta)+1)}=2\sqrt{sec^2(\theta)}=2\sec(\theta)\] and the integral becomes \[\int\frac{\sec^2(\theta)}{\sec(\theta)}d\theta=\int\sec(\theta)d\theta\]

OpenStudy (anonymous):

oh doh not i am not, the two cancels nvm

OpenStudy (anonymous):

thanks satellite

OpenStudy (callisto):

\[\int \frac{1}{\sqrt{t^2-6t+13}}dt\]\[=\int \frac{1}{\sqrt{t^2-6t+3^3-3^2+13}}dt\]\[=\int \frac{1}{\sqrt{(t-3)^2+4}}dt\] Let \(t-3 = 2tan\theta\) \(dt = 2sec^2\theta d\theta\) \[=\int \frac{1}{\sqrt{(2tan\theta)^2+4}}(2sec^2\theta d\theta)\]\[=\int \frac{2sec^2\theta}{\sqrt{(2tan\theta)^2+4}} d\theta\]\[=\int \frac{2sec^2\theta}{\sqrt{(4(tan^2\theta+1)^2}} d\theta\]\[=\int \frac{2sec^2\theta}{\sqrt{(2sec^2\theta)^2}} d\theta\]\[=\int \frac{2sec^2\theta}{2sec\theta} d\theta\]\[=\int sec\theta d\theta\]Amazing :D

OpenStudy (anonymous):

Thanks @Callisto :)